The Apollo 11 spacecraft that landed on the moon in 1969 traveled there at a speed relative to the earth of 1.08 × 104 m/s. To an observer on the earth, how much longer than his own day was a day on the spacecraft?
Answer #1

given v = 1.08*104 mg/s 4-24*60*60 = 86,400 sec t = (to) / ((1-v2 / c2 ))1/2-(86, 400) / ((1-(1.08 * 104 )2 / (3 *108 )2 )-8